OpenStudy (anonymous):

A farmer has three hay fields. The area of one field is 7/8 acre. The are of another field is 5/6 acre. The area of the third field is 11/12 acre. What is the total area of the three fields?

OpenStudy (mathstudent55):

Addition of fractions: write it down and then find a common denominator. \[\frac{ 7 }{ 8 } + \frac{ 5 }{ 6 } + \frac{ 11 }{ 12 } =\]

OpenStudy (anonymous):

Okay what I got was 22 over 24...then what???

OpenStudy (anonymous):

oh, that would be 22 over 26?

OpenStudy (mathstudent55):

What is the least common multiple of 8, 6, and 12. In other words what is the smallest number that is divisible by 8, 6, and 12, in other words, what is the smallest number you can divide 8, 6, and 12 into and get no remainder?

OpenStudy (mathstudent55):

24 is correct. No change every fractio so it has denominator 24

OpenStudy (anonymous):

But nothing goes evenly into 22 evenly except for 11

OpenStudy (anonymous):

So the answer would be 11/12???

OpenStudy (mathstudent55):

\[\frac{ 7 }{ 8 } \times \frac{ 3 }{ 3 } + \frac{ 5 }{ 6 } \times \frac{ 4 }{ 4 } + \frac{ 11 }{ 12 } \times \frac{ 2 }{ 2 }\]

OpenStudy (mathstudent55):

When you multiply each two fractions, you get the same denominator\[\frac{ 21 }{ 24 } + \frac{ 20 }{ 24 } + \frac{ 22}{ 24 }\]

OpenStudy (mathstudent55):

Now to add the fractions, add the numerators and write the same denominator

OpenStudy (anonymous):

2.625 Or 2.6

OpenStudy (mathstudent55):

They add to 63/24 Now you need to reduce

OpenStudy (anonymous):

right...

OpenStudy (mathstudent55):

Divide the numerator and the denominator by 3: 21/8

OpenStudy (mathstudent55):

Now you can write it as a mixed numaral: 2 5/8, or as a decimal 2.625

OpenStudy (anonymous):

HA! Then I got it right (see above). So the answer is 2.6...correct?

OpenStudy (anonymous):

Now how would that be written as a fraction?

OpenStudy (anonymous):

2 5/8 ???

OpenStudy (anonymous):

Okay, I think I got it but not quite sure.

OpenStudy (mathstudent55):

2 5/8 is correct